A straight wire, 0.05 m long, is placed in a uniform magnetic field of magnetic induction 0.04 T. The wire carries a current of 7.5 A, and makes an angle of 60° with the direction of the magnetic field. The arrow on the wire shows the direction of the electron current in the wire. (a) Calculate the magnitude of the force exerted on the wire. (b) Draw a sketch of the wire in the magnetic field and show the direction of the force. (c) Describe the conditions for this force to be a maximum. 


1
Expert's answer
2022-01-12T08:31:52-0500

"F=IBl\sin \alpha=7.5\cdot0.04\cdot 0.05\cdot 0.87=13~mN."

"F= F_{max}\to\sin\alpha=1,"

"F_{max}=IBl=7.5\cdot 0.04\cdot 0.05=15~mN."